A solid disk rotates in the horizontal plane at an angular velocity of 0.038 rad/s with respect to an axis perpendicular to the disk at its center. The moment of inertia of the disk is 0.12 kg · m2. From above, sand is dropped straight down onto this rotating disk, so that a thin uniform ring of sand is formed at a distance of 0.40 m from the axis. The sand in the ring has a mass of 0.50 kg. After all the sand is in place, what is the angular velocity of the disk?

rad/s
Answer #1

Use the law of conservation of angular momentum:


I * ω = ( I + I') * ω'

0.12 * 0.038 = ( 0.12 + (0.50 * 0.40^2)) * ω'

0.00456 = 0.1208ω'

-->ω' = 0.03774834437 rad/s